OpenStudy (anonymous):

Solve the following equation. 5(x - 4) = 12 x=

OpenStudy (potatoes.ramu):

5x-20=12

OpenStudy (potatoes.ramu):

5x=32

OpenStudy (potatoes.ramu):

x=32/5

OpenStudy (anonymous):

5x-20=12 12+20=32 32/5= 6.4